OpenStudy (anonymous):

Jack wants to find the solution to a system of equations using y = x + 3 and y = x^2 + 6x - 21. Betty says that Jack can solve x^2 + 5x - 24 = 0 to find the x-coordinates of the solutions of the system. Explain and demonstrate why Betty is correct.

OpenStudy (anonymous):

y = x + 3 y = x^2 + 6x - 21 When you combine them you get: x + 3 = y = x^2 + 6x - 21 x + 3 = x^2 + 6x - 21 Subtract x +3 from both sides: 0 = x^2 + 5x - 24

OpenStudy (anonymous):

To solve y = x + 3 simultaneously with y = x^2 + 6x - 21, we put them equal to each other to get x^2 + 6x - 21 = x + 3. Putting all the terms on the left gives x^2 + 6x - 21 - x - 3 = 0 which gives x^2 + 5x - 24 = 0. Betty is correct!